### 23 Jan 2024 v1

# Neutron Scattering Studies on the High- Superconductor LaNiO at Ambient Pressure

Abstract
After several decades of studies of high-temperature superconductivity, there is no compelling theory for the mechanism yet; however, the spin fluctuations have been widely believed to play a crucial role in forming the superconducting Cooper pairs. The recent discovery of high-temperature superconductivity near 80 K in the bilayer nickelate LaNiO under pressure provides a new platform to elucidate the origins of high-temperature superconductivity. We perform elastic and inelastic neutron scattering studies on a polycrystalline sample of LaNiO at ambient pressure. No magnetic order can be identified down to 10 K. However, we observe a weak flat spin-fluctuation signal at 45 meV in the inelastic scattering spectra. The observed spin excitations could be interpreted as a result of strong interlayer and weak intralayer magnetic couplings for a double-stripe antiferromagnetic order. The absence of long-range magnetic order in neutron diffraction measurements may be ascribed to a short-range magnetic correlation length perpendicular to the stripe direction. Our results provide crucial information on the spin dynamics and are thus important for understanding the superconductivity in LaNiO.
Introduction. The mechanism of high-temperature (high-) superconductivity is a long-lasting mystery although tremendous progress has been achieved since the discovery of superconductivity in cuprates and iron-based compounds. High- superconductivity in copper-oxide and iron-based superconductors always appears after the suppression of the antiferromagnetic (AFM) order in the parent compounds by chemical doping or pressure and evolves into a superconducting dome with the highest around the critical doping/pressure where the AFM order disappears Scalapino (2012); Tranquada et al. (2014); Keimer et al. (2015); Dai (2015). As the AFM order is suppressed by chemical doping, the spin fluctuations persist throughout the phase diagram and form a neutron spin resonant mode below at certain energies around the AFM wave vectors of the parent compounds. The temperature dependence of the intensity of the neutron spin resonant mode behaves like a superconducting order parameter and the resonant energy can be linearly scaled with the corresponding and superconducting gap Eschrig (2006); Sidis et al. (2007); Christianson et al. (2008a); Yu et al. (2009). All these characteristics suggest that magnetism is crucial to the mechanism of the high- superconductivity, and the AFM spin fluctuations could be the “glue” for the formation of superconducting Cooper pairs.
Seeking new unconventional high- superconductors is the key to understanding the high- mechanism and extending the application of superconductivity. Nickelates have been proposed to be a promising material platform to search for high- superconductors due to the similarity of the lattice and electronic structure with those of the cuprates Anisimov et al. (1999). Among different families of nickelates, NiO ( = La, Nd, Pr) were predicted to be the most probable to achieve superconductivity due to the same electronic configuration of Ni (3) with Cu in the cuprates. Superconductivity was indeed observed in NdSrNiO thin films ( 9-15 K) in 2019 and later confirmed in other thin film samples with similar compositions Li et al. (2019); Osada et al. (2020); Wang et al. (2022); Gu and Wen (2022). Additionally, it is interesting to note that spin-wave-like magnetic fluctuations of NdSrNiO from resonant inelastic X-ray scattering (RIXS) measurements give comparable magnetic exchange couplings to that in copper-oxide superconductors Lu et al. (2021); Coldea et al. (2001).
Recently, superconductivity with up to 80 K was reported in single-crystal samples of the Ruddlesden-Popper phase bilayer nickelate LaNiO [Fig. 1(a)] with pressure above 14 GPa Sun et al. (2023). Either zero-resistance or a diamagnetic response corresponding to the superconducting transition in LaNiO has been confirmed on single-crystal or polycrystalline samples in subsequent reports Hou et al. (2023); Zhang et al. (2023a); Wang et al. (2023); Zhou et al. (2023). X-ray photoemission spectroscopy measurements indicated the valence state of the Ni ions in LaNiO is a mixture of Ni and Ni, which may result in a charge density wave (CDW) and a spin density wave (SDW)Liu et al. (2023a); Taniguchi et al. (1995). Electrical transport and specific heat measurements on single crystal samples indeed revealed transition-like anomalies in resistance and heat capacity at 153 and 110 K Liu et al. (2023a). An optical study observed a CDW-like transition at 115 K Liu et al. (2023b). A recent SR study provided evidence of the existence of the SDW in a LaNiO powder sample at 150 K Chen et al. (2023). However, neutron diffraction measurements did not observe long-range magnetic order in powder samples Ling et al. (1999). Thus, an elucidation of the nature of the magnetic ground state, the properties of the spin fluctuations, and the magnitudes of the magnetic exchange interactions of LaNiO are crucial to understanding the mechanism of the high- superconductivity.
Although the angle of the bilayer Ni-O-Ni along the axis evolves from 168 to 180 under pressure Sun et al. (2023), the spin fluctuations and magnetic exchange interactions under ambient conditions should be comparable to those at high pressure. In this paper, we report neutron diffraction and systematic inelastic neutron scattering (INS) measurements on a polycrystalline LaNiO sample at ambient pressure. No long-range magnetic order can be identified at 10 K. The INS data reveal a non-dispersive spin excitation signal at meV, which has been used to estimate the magnitude of the magnetic exchange couplings. We find that an alternate double stripe AFM order with only short-range correlations or very small magnetic moments exhibiting a strong interlayer coupling and weak intralayer couplings matches the experimental observations. In addition, dispersive and non-dispersive phonon modes are observed. By employing density functional theory (DFT) calculations, we have computed the dispersion and the density of states (DOS) of the phonon modes, which are roughly consistent with the INS spectra. Our results reveal that the strong interlayer magnetic exchange couplings dominate the spin fluctuations, which is distinct from the situation for the cuprates and iron-based superconductors.
Characterizations and experimental details. Figure 1(c) shows basic bulk characterization of the sample. The temperature dependence of the resistivity goes up slightly at low temperatures. The magnetization at = 1000 Oe also increases towards lower temperatures. Both upward trends at low temperatures suggest the existence of oxygen vacancies Zhang et al. (1994). Neutron diffraction experiments were performed on the general purpose powder diffractometer (GPPD) Chen et al. (2018) at China Spallation Neutron Source (CSNS) at 10 and 160 K. The data were refined with the Rietveld method using the GSAS software package Toby (2001). The low-energy INS spectra at 2 and 200 K were collected on the time-of-flight cold neutron spectrometer PELICAN Yu et al. (2013) at the Australian Centre for Neutron Scattering (ACNS) at the Australian Nuclear Science and Technology Organisation (ANSTO) with an incident neutron energy = 3.7 meV. The high-energy spectra were obtained from the time-of-flight spectrometers MERLIN Bewley et al. (2006) and MAPS Ewings et al. (2019) at the ISIS Neutron and Muon Source at UK with s = 15, 22.6, 36, 66.3, 150.6, and 500 meV (MAPS) at 7 and 300 K. We used 5 g powder samples of LaNiO synthesized by the solid-solid reaction method Liu et al. (2023a) for the experiments on GPPD and PELICAN. 20 g powder samples synthesized in the same way were used for the experiments on MERLIN and MAPS. The INS data from PELICAN were analyzed using the LAMP software package. The data collected at MERLIN and MAPS were reduced and analyzed using the software packages MANTIDPLOT Arnold et al. (2014) and DAVE Azuah et al. (2009). The scattering intensities from aluminum have been removed by subtracting the data from an empty can for the PELICAN and MERLIN experiments. The spin excitation simulations were performed using the SpinW software package Toth and Lake (2015). The phonon dispersions and DOS were calculated using the finite displacement method as implemented in the Phonopy Togo and Tanaka (2015) software in combination with the Vienna Simulation Package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996).
Neutron diffraction results. Figure 1(d) presents the neutron powder diffraction data measured at 10 K. No extra peaks can be seen when comparing the data between 10 and 160 K (see supplementary information (SI) SI for details). Our refinements reveal two structures of LaNiO with space groups of (No. 63) and (No. 69). The refined compositions at 10 K for the two phases are 56.1% LaNiO and 41.0% LaNiO for the and structures, respectively SI . A fraction of 2.9% LaNiO impurity is also identified. Based on the refined results, there are more vacancies on the inner apical oxygen sites for the phase, consistent with a recent transmission electron microscopy experiment Dong et al. (2023). We thus use LaNiO to refer to our sample hereafter. It should be noted that LaNiO with the structure has also been reported at ambient pressure Zhang et al. (1994).
Inelastic neutron scattering results. Figures 2(a)-2(d) present the INS spectra measured with different s at 2 K on PELICAN and 7 K on MERLIN and MAPS. Identical measurements at 200 and 300 K are presented in the SI SI . One cannot see obvious spin fluctuations from the color maps at low . At higher and energy ranges weak acoustic phonon modes in Fig. 2(b) and flat optical phonon modes in Fig. 2(c) can be observed. To inspect the INS spectra more closely, we performed one-dimensional (1D) cuts along the and energy axes from both the low-temperature and high-temperature data. In the low range of [0.6, 1.9] Å as shown in Fig. 2(e), where the intensity of phonons should be weak, the difference in intensities around 1 meV may be ascribed to temperature-enhanced spin fluctuations followed by the Bose factor. In Fig. 2(f), we present 1D cuts along the axis by integrating the energies between 2.5 and 3.5 meV from the map in Fig. 2(b). The cut at 300 K in the lower panel of Fig. 2(f) reveals obvious acoustic phonons, whose intensities become much weaker at 7 K. Interestingly, obvious signals above the background at 7 K can be observed below 3 Å as shown in the upper panel of Fig. 2(f). The intensities observed at low and low temperature of 7 K should be magnetic in origin.
Since weak magnetic excitations at higher energies may be embedded in the INS spectra, we performed direct data subtractions between 7 and 300 K, as shown in Fig. 2 and the SI SI . The subtraction results are shown in Fig. 3. Phonon intensities at 300 K are expected to be much stronger than at 7 K. However, a flat mode centered at 45 meV appears for the data collected at MERLIN [Figs. 3(a) and 3(b)] and MAPS [Fig. 3(c)]. The intensities decrease as increases as shown in the inset of Fig. 3(a)], indicating the magnetic origin. We show a 1D cut in the range of [1.8, 4] Å along the energy axis to see the exact energy of the magnetic excitations in Fig. 3(d). The peak of the magnetic mode centers at 45 meV. The tiny positive intensities around meV are likely to originate from the mismatch of the line width of the optical phonon modes between 7 and 300 K. The dependence of these intensities does not follow a magnetic form factor. Thus, we only consider the features around 45 meV to be magnetic excitations.
We now attempt to estimate the magnetic exchange couplings by assuming a certain magnetic structure and comparing its magnetic excitation spectrum to the INS data. Several types of magnetic orders have been tested and a range of possible combinations of exchange couplings can be found in the SI SI . The following information was taken to determine the magnetic order and set of s: (1) The signature of SDW revealed by SR Chen et al. (2023) and the absence of magnetic Bragg peaks in neutron diffraction, suggesting that the magnetic order is short-range or involves only small magnetic moments. (2) An AFM magnetic structure with alternate stripes of magnetic and non-magnetic Ni atoms is highly possible, given the short-range character and the mixture of Ni and Ni in LaNiO with the same proportion Liu et al. (2023a). (3) The intensities of the dispersive magnetic excitations at low and low energies are weak, while the flat magnetic excitation at 45 meV is prominent. (4) The interlayer coupling should be much stronger than the intralayer couplings due to the half-filling of the 3 orbital and the quarter-filling of the 3 orbitals Sun et al. (2023); Yang et al. (2023a).
Figure 3(e) shows a double stripe-type AFM order with magnetic and non-magnetic Ni atoms. We employ a Heisenberg Hamiltonian by considering the effective intralayer exchange couplings , , , and the inter-layer coupling . In Fig. 3(f), we present the calculated magnetic excitation spectrum using = 5, = -2, = 1, and = 36 meV with the magnetic order shown in Fig. 3(e), which matches the INS data well. The double stripe-type AFM order in Fig. 3(e) and the set of magnetic exchange couplings obey the above constraints. We note the observed INS spectrum cannot give a unique magnetic order. However, the relative strong interlayer coupling and weak intralayer couplings should be reliable.
The acoustic phonon branches have been shown in Figs. 2(b) and 2(f). The 1D constant cuts focusing on the optical phonons are exhibited in Figs. 4(a) ( = 66.3 meV) and 4(b) ( = 160 meV). The most notable features in Fig. 4(a) are the strong peaks around 11 and 23 meV and weak peaks around 30, 34, and 40 meV, which all have stronger intensities at 300 K than that at 7 K, confirming their phononic origin. Similarly, the INS peaks around 58, 68, and 78 meV that correspond to the flat bands in Figs. 2(c) can be identified [Fig. 4(b)]. The phonon calculations were conducted using the DFT method. The dispersion of the phonon modes along the high-symmetry directions is shown in Fig. 4(c). A large number of phonon branches with different levels of dispersion can be seen. The vibrational density of states is presented in Fig. 4(d), where the peak features around 10-15, 30, 40, 55, and 70 meV can be recognized. These peaks are roughly consistent with the experimental INS peaks in Figs. 4(a) and 4(b). The calculated partial DOS shows that the higher flat bands are dominated by the vibration of oxygen atoms as expected.
Discussion and Conclusion. Our detailed comparisons of the INS spectra at low and high temperatures reveal convincing evidence of magnetic excitations in LaNiO. Although the results are obtained at ambient pressure, the proposed magnetic exchange couplings are important for understanding the mechanism of superconductivity and the pairing gap symmetry in the superconducting state under high pressure Yang et al. (2023b); Zhang et al. (2023b); Liu et al. (2023c); Qu et al. (2023). The results also indicate the existence of spin fluctuations is universal among copper-oxide, iron-based superconductors, and nickelate high superconductors.
We briefly discuss the reasons for the weakness of the spin fluctuations in LaNiO. First of all, we should note that LaNiO is a metallic system that shows itinerant characteristics. Some of the nickel ions may remain nonmagnetic or in low spin states. The possible alternate double stripe AFM order with a short-range correlation perpendicular to the stripe direction will further suppress the long-range magnetic order and weaken the intensity of magnetic excitations. Secondly, as shown in Fig. 1(b), the 3 electrons can form -bonding and anti-bonding bands within a bilayer unit and the 3 electrons hybridize with the 3 electrons within one Ni-O layer, which give rise to nearly half-filled 3 orbit and quarter-filled 3 orbit. The quarter-filled character of the 3 orbit represents a weak intralayer exchange coupling which further reduces the intensity of spin fluctuations. This behavior is different from that in cuprate and iron-based superconductors, where the dominant intralayer exchange couplings lead to the formation of in-plane long-range magnetic order/correlation. Thirdly, polycrystalline samples always give weak signal intensity in INS spectra due to the averaging effect when compared with the results from single-crystal samples with a similar mass. In powder samples of iron-based superconductors, the spin fluctuations are almost invisible Qiu et al. (2008); Christianson et al. (2008b); Mittal et al. (2008), except for the case when there is a strong neutron spin resonance in superconducting samples Christianson et al. (2008a).
In conclusion, we report the elastic and inelastic neutron scattering results of a polycrystalline LaNiO sample at ambient pressure. No long-range magnetic order can be identified at 10 K. Our INS results reveal magnetic excitations in LaNiO, indicating that the nickelate superconductors might be in a similar position to the copper-oxide and iron-based superconductors when considering the role of magnetism in the mechanism for high- superconductivity. Estimating the magnitude of the magnetic exchange couplings by assuming an alternate double stripe AFM order results in a strong interlayer coupling and weak intralayer couplings. Our work provides crucial information to the community for determining a realistic mechanism of superconductivity in LaNiO.
Acknowledgements. We thank Prof. D. L. Feng for sharing the RIXS data in the ‘Workshop on High- Nickelate Superconductors’. Work at SYSU was supported by the National Key Research and Development Program of China (Grant Nos. 2023YFA1406500, 2023YFA1406002), the National Natural Science Foundation of China (Grant nos. 12304187, 12174454, U21301001), the Guangdong Basic and Applied Basic Research Funds (grant no. 2021B1515120015), Guangzhou Basic and Applied Basic Research Funds (grant nos. 202201011123, 2024A04J6417), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (grant no. 2022B1212010008), and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (Grant No. 23qnpy57). This work is based on experiments (Proposal P15658) performed at Australian Centre for Neutron Scattering (ACNS) at Australian Nuclear Science and Technology Organisation (ANSTO), Australia. The neutron experiment at ISIS Neutron and Muon Source in UK was performed under a user program (Proposal RB2300045 and RB2300079). The neutron diffraction experiment at CSNS in China was performed under the proposal P1822113000042.
Supplementary Information: Neutron Scattering Studies on the High- Superconductor LaNiO at Ambient Pressure
I Neutron Scattering Results
Here we present the neutron scattering raw data that are referred to in the main text. We first compare the neutron powder diffraction results collected on GPPD at 10 K and 160 K in Fig. 5. No new peaks or clear intensity gain on the nuclear peaks from magnetic order can be identified at 10 K. Structural models are refined against the neutron diffraction data with the Rietveld method using the GSAS software package. The refined results are shown in Fig. 6, and the refined structure parameters at 10 and 160 K are summarized in Tables 1, 2, 3, 4.
In Fig. 7, the corresponding high-temperature INS spectra to those shown in Fig. 2 of the main text are presented. These results were subtracted from the corresponding low-temperature data in Fig. 2 to give the difference shown in Fig. 3 of the main text. In Fig. 8, the raw data used for the 1D INS spectra in Fig. 4(a) are presented.
| Lattice | (No. 63) | = 5.3956(4)Å | = 5.4491(4)Å | = 20.598(3)Å |
| atom | occupancy | |||
| La1 | 0.2500 | 0.2453 | 0.3031 | 1 |
| La2 | 0.2500 | 0.2362 | 0.5000 | 1 |
| Ni | 0.2500 | 0.2509 | 0.1059 | 1 |
| O1 | 0.5000 | 0.0000 | 0.0861 | 1 |
| O2 | 0.2500 | 0.2126 | 0.1996 | 1 |
| O3 | 0.0000 | 0.5000 | 0.1055 | 1 |
| O4 | 0.2500 | 0.3014 | 0.0000 | 0.79(6) |
| Lattice | (No. 69) | = 5.3816(2)Å | = 5.4538(3)Å | = 20.470(2)Å |
| atom | occupancy | |||
| La1 | 0.0000 | 0.0000 | 0.5000 | 1 |
| La2 | 0.0000 | 0.0000 | 0.3178 | 1 |
| Ni | 0.0000 | 0.0000 | 0.0942 | 1 |
| O1 | 0.0000 | 0.0000 | 0.0000 | 0.98(8) |
| O2 | 0.0000 | 0.0000 | 0.2064 | 1 |
| O3 | -0.2500 | 0.2500 | 0.0934 | 0.98(2) |
| Lattice | (No. 63) | = 5.4015(4)Å | = 5.4491(4)Å | = 20.616(3)Å |
| atom | occupancy | |||
| La1 | 0.2500 | 0.2453 | 0.3031 | 1 |
| La2 | 0.2500 | 0.2362 | 0.5000 | 1 |
| Ni | 0.2500 | 0.2509 | 0.1059 | 1 |
| O1 | 0.5000 | 0.0000 | 0.0861 | 1 |
| O2 | 0.2500 | 0.2126 | 0.1996 | 1 |
| O3 | 0.0000 | 0.5000 | 0.1055 | 1 |
| O4 | 0.2500 | 0.3014 | 0.0000 | 0.81(6) |
| Lattice | (No. 69) | = 5.3880(2)Å | = 5.4531(3)Å | = 20.489(2)Å |
| atom | occupancy | |||
| La1 | 0.0000 | 0.0000 | 0.5000 | 1 |
| La2 | 0.0000 | 0.0000 | 0.3178 | 1 |
| Ni | 0.0000 | 0.0000 | 0.0942 | 1 |
| O1 | 0.0000 | 0.0000 | 0.0000 | 0.98(8) |
| O2 | 0.0000 | 0.0000 | 0.2064 | 1 |
| O3 | -0.2500 | 0.2500 | 0.0934 | 0.99(2) |
II Linear Spin-Wave Theory Simulation
Since only weak and non-dispersive magnetic fluctuations can be observed in our data, we cannot do quantitative fitting to the data but perform qualitative estimations based on the results. Our strategy to simulate the spin excitations is to assume a magnetic structure and suitable exchange couplings to produce a similar excitation spectrum in the experiment. We note the linear spin-wave theory (LSWT) is suitable for long-range ordered magnetic systems with localized moments. Our sample shows no long-range magnetic order. The assumed magnetic structure is only used for LSWT simulation in SpinW software package.
Several magnetic structures with different combinations of exchange couplings are compared below. We first exclude the well-known Néel-type and A-type antiferromagnetic orders [Fig. 9], since the calculated powder-averaged INS spectra give very different features from our experimental results.
We considered two types of stripe AFM orders with the stripes along the -axis of the crystal. The first type of stripe AFM order is alternate single stripes of magnetic and non-magnetic Ni atoms [Fig. 9(c)], named single stripe-type AFM, and is excluded due to the deviation from the experimental results. The second stripe AFM order is the alternate double stripes of magnetic and non-magnetic Ni atoms [Fig. 10(a)], named double stripe-type AFM, and we find a range of exchange parameters that can reproduce the experimental spectra.
The weak and non-dispersive magnetic fluctuation can only be observed in the subtracted data at high energy, which might indicate the spin excitations are rather weak at low-energy regions. This feature reminds us that one cannot only consider the nearest neighbor (NN) as shown in Fig. 10(b), since it can generate strong intensities in the low-energy region. The and are necessary to push the spectral weight from the low-energy region to the high-energy region as shown in Figs. 10(c)-(e). Thus, we focused on the combination of , , , and to estimate the range of the couplings. Since LaNiO is an itinerant system, which means that there may be no local moments, we directly use the effective , , , , and below. We fix the exchange couplings that are further than and only discuss the relation between and . Figure 11 presents the LSWT simulations of the spectra with a wide range of the parameters of and (from dominant to dominant ). One can see the dominant cases in Figs. 11(d)-(i) can describe the experimental results better than the dominant cases. Interestingly, the combinations of and that can well reproduce the experimental spectra follow a linear relation of = 5 + 61 as that shown in Fig. 12. At last, we estimate the parameters of and that can reproduce the spectra similar to the experimental results and present them in Fig. 12 between the two dashed lines.